OpenStudy (anonymous):

Solve: 5x − 10 + 12 = 22

OpenStudy (lgbasallote):

first, combine like terms (in this case -10 and 12) \[\implies 5x + 2 = 22\] now subtract 2 from both sides \[\implies 5x + 2 - 2 = 22 - 2\] combine like terms \[\implies 5x = 20\] do you know the rest?

OpenStudy (anonymous):

5x=20 x=4

OpenStudy (lgbasallote):

or do you need more help?

OpenStudy (anonymous):

5x-10=22-12 5x=10+10 5x/5=20/5 x=4

OpenStudy (anonymous):

yeah i got it thanks @lgbasallote